OpenStudy (agl202):

I think it's either C. or B. Rewrite f(x) = −2(x − 3)2 + 2 from vertex form to standard form. f(x) = −2x2−18 f(x) = −2x2 + 12x + 20 f(x) = −2x2 + 12x − 16 f(x) = 4x2− 24x + 38

OpenStudy (jango_in_dtown):

−2(x − 3)^2 + 2=-2(X^2-6X+9)+2=-2x^2+12x-18+2 =-2x^2+12x-16

OpenStudy (agl202):

C? I was thinking mostly C. Thx... ;)

OpenStudy (agl202):

It was right! :) <3